# Constrained and Ordered Level Planarity Parameterized by the Number of Levels

### v1

@@ 4 Tractability of 3-Level Constrained Level Planarity
In this section, we sketch our polynomial-time algorithm for 3-level CLP (see Theorem 1.2). For simplicity, we assume that the given level graph $\mathcal{G}$ is constrained level planar and show how to compute a constrained level planar drawing of $\mathcal{G}$. We ask the reader to consult the figures to which we refer in order to get an intuitive understanding of the notions that we cannot define formally due to space constraints. For details and proofs, see Appendix E.
Without loss of generality, we make the following assumptions: (i) $\mathcal{G}$ is proper; otherwise we subdivide each edge that crosses level 2. (ii) $\mathcal{G}$ has no isolated vertices; they can easily be inserted in a postprocessing step. (iii) The component–constraint graph $H$ is strongly connected. This directed graph has a node for each connected component (for short, component) of $\mathcal{G}$ and an arc for each pair $(C,C^{\prime})$ of components such that there is a constraint from a vertex in $C$ to a vertex in $C^{\prime}$.
@@
We often refer to level 1 as bottom level, to level 2 as middle level, and to level 3 as top level. Similarly, we call the pair of bottom and middle level lower band and the pair of middle and top level upper band. Our algorithm successively adds new constraints to the given constrained level graph $\mathcal{G}$ that do not violate planarity. We deduce these new constraints from the structure of $(G,\gamma)$ and from the current set of constraints. In the end, this yields a total order of the vertices for each of the three levels and, hence, a constrained level planar drawing of $\mathcal{G}$. In the very beginning and whenever we add new constraints, we exhaustively add the following implicit constraints:
  • transitivity: $\forall a,b,c\in V(G),\forall i\in[3]\colon a\prec_{i}b\land b\prec_{i}c\Rightarrow a\prec_{i}c$
  • planarity: $\forall ab,cd\in E(G)\text{ with }i:=\gamma(a)=\gamma(c)\text{ and }j:=\gamma(b)=\gamma(d)\colon\\ a\prec_{i}c\Rightarrow(b\prec_{j}d\lor b=d)$

### v2

@@ 6 Tractability of 3-Level Constrained Level Planarity
In this section we show that, given a constrained level graph $\mathcal{G}=(G,\gamma,(\prec_{i})_{i\in[3]})$ of height 3, we can decide in polynomial time whether $\mathcal{G}$ admits a constrained level planar drawing. We assume that $\mathcal{G}$ is level planar (which can be tested in linear time [22, 25]) and proper (otherwise we subdivide edges that connect levels 1 and 3). Besides referring to levels 1, 2, and 3 by their indices, we also call them bottom level, middle level, and top level, respectively. Furthermore, we call the pair of bottom and middle level lower band and the pair of middle and top level upper band.
@@
Throughout this section, we successively add new constraints to the given constrained level graph $\mathcal{G}$ that we deduce from the structure of $G$, $\gamma$, and the current set of constraints. In the end, this yields a total order of the vertices for each of the three levels that corresponds to a constrained level planar drawing of $\mathcal{G}$, or we come to the conclusion that $\mathcal{G}$ does not admit such a drawing. In the very beginning and whenever we add new constraints, we exhaustively add the following implicit constraints:
  • Transitivity: For every triplet $(a,b,c)$ of vertices on the same level, if there is a constraint from $a$ to $b$ and from $b$ to $c$, then there is also a constraint from $a$ to $c$. Formally,
    $\forall a,b,c\in V(G)$ with $i:=\gamma(a)=\gamma(b)=\gamma(c)\colon(a\prec_{i}b)\land(b\prec_{i}c)\Rightarrow(a\prec_{i}c)$.
  • Planarity: For every pair $(ab,cd)$ of edges such that $a$ and $c$, and $b$ and $d$ lie on the same level, if there is a constraint from $a$ to $c$, then there is also a constraint from $b$ to $d$ or $b=d$. Formally, $\forall ab,cd\in E(G)\text{ with }i:=\gamma(a)=\gamma(c)\text{ and }j:=\gamma(b)=\gamma(d)\colon$ $(a\prec_{i}c)\Rightarrow(b\prec_{j}d)\lor(b=d)$.